### v2

# thanks: A footnote to the article title

Abstract
As the workhorse of electronic structure methods, density functional theory (DFT) has become essential for molecule and materials discovery, deploying databases covering vast chemical and composition spaces, predicting chemical and electrochemical reactions, and developing machine learning potentials. The extensive application of DFT in chemistry and materials science requires predictions achieving ”chemical accuracy,” a goal hindered by the unknown functional form for the exact exchange and correlation (XC) energy. A recent meta-generalized gradient approximation, the restored-regularized strongly constrained and appropriately normed (r2SCAN) XC functional, fulfils 17 exact constraints of the XC energy, and has significantly boosted the accuracy of predictions for molecules and materials. However, r2SCAN still appears inadequate at predicting material properties of open d and f transition-metal strongly correlated compounds, including band gaps, magnetic moments, and oxidation energies. Prediction inaccuracies of r2SCAN arise from functional and density-driven errors, mainly resulting from the DFT self-interaction error. Here, we propose a novel method termed r2SCANY@r2SCANX to mitigate the pernicious self-interaction error of XC functionals for the accurate simulations of electronic, magnetic, and thermochemical properties of transition metal oxides. r2SCANY@r2SCANX utilizes different fractions of exact Hartree-Fock exchange: X to define the electronic density, and Y to set the density functional approximation for the energy. We show that r2SCANY@r2SCANX simultaneously addresses functional-driven and density-driven inaccuracies, mitigating the self-interaction error in DFT. Building just on 1 (or maximum 2) universal parameters, we demonstrate that r2SCANY@r2SCANX improves upon the r2SCAN predictions for 18 highly correlated oxides and even outperforms the highly parameterized DFT(r2SCAN)+U method –the state-of-the-art approach to predict strongly correlated materials. We demonstrate that the typical O2 overbinding inaccuracies of the local density and the gradient XC approximations, with errors ranging between –2.2 and –1.0 eV/O2, reduce to –0.3 eV/O2 for r2SCAN, and further reduce to less than $\sim$0.03 eV/O2 using any combination of X in r2SCAN10@r2SCANX, including X=0. Prediction uncertainties for oxidation energies and magnetic moments of transition metal oxides are significantly reduced by r2SCAN10@r2SCAN50 and band gaps with r2SCAN10@r2SCAN. r2SCAN10@r2SCAN50 diminishes the density-driven error present in r2SCAN and r2SCAN10. We demonstrate that the computationally efficient r2SCAN10@r2SCAN is nearly as accurate as the global hybrid r2SCAN10 for oxidation energies. This suggests that accurate energy differences can be achieved through rate-limiting self-consistent iterations and geometry optimizations using the efficient r2SCAN. Subsequently, a more expensive nonlocal functional, such as a hybrid or self-interaction correction, can be applied in a fast-to-execute single post-self-consistent calculation, as in r2SCAN10@r2SCAN. Compared to computationally demanding r2SCAN10, r2SCAN10@r2SCAN appears 10 to 300 times faster.
I Introduction
Electronic structure methods, especially density functional theory (DFT),1 have become essential for materials discovery, enabling the prediction of properties and behaviors of technologically relevant materials.2, 3, 4, 5, 6 This has resulted in the widespread use of DFT for extensive materials databases, such as the Alexandria Materials Database,7 AFLOW,8 GNoME,9 the Materials Project,3 the NREL materials databases,10 and OQMD.2 These vast databases provide a critical foundation for materials science research, enabling direct comparison of computed properties with experimental data, developing predictive models, and lately, foundational training sets for machine learning potentials,11, 12, 13 thus necessitating highly accurate datasets approaching “chemical accuracy”.
Nevertheless, DFT in the Kohn-Sham formulation requires approximations for the unknown exchange and correlation (XC) functional, which has been a matter of intense research for the past 60 years.14, 15, 16 Among the density functional approximations (DFAs), the local spin density approximation (LSDA),17 and the generalized gradient approximation (GGA)18 exhibit significant inaccuracies in predicting structural parameters, energetics, and electronic properties, especially for strongly correlated systems. DFA and XC are synonyms and will be used interchangeably.
In this context, the strongly constrained and appropriately normed (SCAN)19 and its regularized and computationally efficient version, r2SCAN,20 are highly accurate meta-GGA functionals. Both SCAN and r2SCAN satisfy 17 known exact constraints in constructing XC functionals, ensuring a well-balanced description of XC effects for a wide range of systems.21 Another constraint-based meta-GGA is LAK, which accurately describes the electronic bonding (at least for main group semiconductors) and band gaps of molecules and solids.22 SCAN and r2SCAN have done admirably well in improving the quality of predictions of molecules and materials relative to LSDA and GGA.23, 24, 25, 26, 27, 28 Note that most databases discussed above rely entirely on the predictive capabilities of GGA, and ad hoc corrections, including GGA+U,29, 30 and more recently r2SCAN.31
Challenges persist when SCAN and r2SCAN are used to predict several valuable material properties of open d and f transition-metal compounds, including band gaps, magnetic moments, and oxidation (reduction) energies. While SCAN and r2SCAN XC functionals reduce the self-interaction error (SIE) magnitude compared to standard semi-local GGA XC functionals, this pernicious error remains.23, 24, 32, 33, 34, 35, 36, 20
Pragmatic but material- and property-dependent solutions to address the SIE in GGA and meta-GGA functionals involve parametrizing ad hoc on-site Hubbard U corrections, such as GGA(LSDA)+U,37 and r2SCAN(SCAN)+U.25, 26, 27, 29, 30, 38, 39, 40 Although numerically accurate and computationally affordable, these XC functionals+U approaches generally lack transferability as the U parameters depend on material chemistry, dimensionality, and the oxidation states of transition metals. Numerous strategies exist to fit the U values to the thermochemical data of redox reactions,25, 26, 27, 29, 30, 39, 40 band gaps.37, 41 U values can also be derived from linear response,42, 43 or even with machine learning models.44, 45 A persistent issue exists: when a U-specific value is applied to oxidation energies, it is not evident that this same numerical value of U will accurately replicate the band gaps (or magnetic moments) under similar conditions in materials.29, 30
A more universal strategy for addressing SIE in materials is that of hybrid XC functionals,46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58 where a portion of the (LSDA, GGA, or meta-GGA) exchange term is replaced by a fraction of exact Hartree-Fock (HF) exchange. Global and range-separated hybrids have been successfully applied to transition-metal oxides (MiOjs).46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64 Although in most hybrid functionals the amount of HF exchange remains a parameter set a priori (as in DFT+U),65, 66 hybrid functionals appear less sensitive to material types or particular reactions and properties.46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 67 A few non-empirical approaches to determine the amount of HF exchange also exist.68, 69, 70, 71, 72, 73, 74, 75, 63 In the same vein, this paper explores several global r2SCAN-based hybrid functionals with Xof exact HF exchange, termed r2SCANX.
Building on the idea that HF provides self-interaction-free, uncorrelated densities, an elegant solution is to use HF electronic charge densities to compute r2SCAN total energies.76, 77, 78, 79, 80, 81, 82 This approach is called the Hartree-Fock density functional theory or DFA@HF.82 r2SCAN@HF has been successfully applied to molecules.76, 77, 78, 79, 80, 81, 82 In almost all cases, r2SCAN@HF is significantly more accurate or slightly less accurate than r2SCAN.82 There are instances where HF densities outperform specific XC functionals when describing molecular systems, especially charge transfer reactions. In other cases, an ”unconventional error cancellation” discussed in Refs. 82, 83, 84 occurs, leading to surprisingly good numerical results, even when HF electronic densities are less accurate than r2SCAN densities. This paper proposes an r2SCAN@HF-like approach to predicting the electronic and magnetic properties of crystalline open-d 1st row transition metal oxides and their oxidation reactions without relying on the unconventional error cancellation, which we will show does not occur for our transition metal oxides.
Unconventional error cancellations observed in r2SCAN@HF is demonstrated in Eq. 1,85, 86, 87, 88 expressing the error of the DFT total energy $E$ as a sum of a functional-driven error, FE –intrinsic to the DFA– and a density-driven error –DE caused by inaccuracies carried by the electronic charge density $n$.
$\Delta E_{DFA}$ $=E_{DFA}\left[n_{DFA}\right]-E_{exact}\left[n_{exact}\right]=\mathrm{FE}+\mathrm{DE},$ (1)
$\mathrm{FE}$ $=E_{DFA}\left[n_{exact}\right]-E_{exact}[n_{exact}],$
$\mathrm{DE}$ $=E_{DFA}\left[n_{DFA}\right]-E_{DFA}[n_{exact}]$
where $n_{DFA}$ is the “inaccurate” charge density provided by the DFA, and $n_{exact}$is the unknown exact density. To evaluate FE and DE of Eq. 1, an exact (or nearly exact) electron density is needed. Refs. 82, 83, 84 utilized the coupled-cluster CCSD(T) density to be exact in a few representative sp-bonded systems, and found for those systems that it gave FE and DE values very close to those that took, as the exact density, the r2SCAN50 density, where r2SCAN50 is the global hybrid of r2SCAN with 50% of HF exact exchange. Then r2SCAN50 was used as the proxy for the exact density in all such systems.82, 83, 84 The r2SCAN50 density was understood to provide correct electron transfers89 from one atomic site to another, a feature of the electron density to which the total energy appears especially sensitive. Increasing the exact-exchange fraction X is expected90 to increase the tendency to put an integer number of electrons on each atom (or ion). It has been our experience with self-consistent DFA that DE is typically much smaller in magnitude than FE in transition-metal oxidation energies, as it is in the barrier heights of main-group molecules and the binding energies of water clusters.82, 83, 91
In sp-bonded systems where electron transfer is absent or very minimal, the r2SCAN densities were found to be much more accurate than the HF densities.92, 93 The improvement in density from HF to LSDA, to PBE, and to r2SCAN shows the predictive power of including systematically more exact constraints in the DFA.21, 83, 84 Refs. 21, 83, 84 concluded that DFA@HF, in which a DFA is evaluated on the HF orbitals and densities, often works through an unconventional error cancellation between the FE of the DFA and the DE of the HF density, $E_{DFA}[n_{HF}$]-$E_{DFA}\left[n_{exact}\right]$. For example, when using DFA@HF, Kanungo et al.,82 observed that unconventional error cancellation favorably improved barrier heights in molecular reactions, which are also impacted mainly by SIE. This observation has been linked to error cancellation between large negative values of FE (Eq. 1) counterbalanced by correspondingly large positive values of $DE$; the latter caused by uncorrelated HF charge densities.82
For transition metal oxides, in this paper, we will show that Hartree-Fock density functional theory fails because there is no such cancellation of errors in these systems. Therefore, more accurate methodologies for the SIE in transition metal oxides will be introduced. By correcting the FE and DE (Eq. 1) independently, here, we address the SIE of XC functional in predicting the electronic, magnetic, and thermodynamic properties of transition metal oxides. This is achieved by mixing different fractions of exact HF exchange in the DFA used to evaluate the total energy and to compute the electronic charge density. This strategy is applied with the previously proposed r2SCAN, resulting in a generalized r2SCANY@r2SCANX approach, where Y is the percentage of exact HF exchange mixed with r2SCAN functional used to evaluate energy. X represents the % of exact HF exchange mixed with r2SCAN used to compute the orbitals, and hence the electronic charge density.
We will demonstrate that when r2SCANY@r2SCANX is applied to transition metal oxides’ oxidation energies, as values of Y and/or X increase, the error in oxidation energies drops noticeably and minimizes at Ysimilar-$\sim$10and Xsimilar-$\sim$50%. We will propose a computationally efficient, non-self-consistent, r2SCAN10@r2SCAN approach that significantly improves the accuracy of oxidation energies compared to r2SCAN and on par with highly parametrized r2SCAN+U approaches. We will also demonstrate that r2SCAN10@r2SCAN outperforms r2SCAN and r2SCAN+U in terms of accuracy for band gaps.
II Results
II.1 Improving Ab initio Predictions of meta-GGA-type r2SCAN Functionals
All XC functionals inherently suffer from functional-driven and density-driven errors (Eq. 1). SIE, a significant part of both errors, leads to inaccurate predictions of energetics in molecules and materials. LSDA, GGA, meta-GGA XC functionals struggle to accurately describe properties, such as reaction energies, inter-atomic charge transfer, and electronic structure, particularly in systems with strongly localized d (or f) electrons, such as MiOjs.10, 25, 26, 29, 30, 37, 38, 39, 39, 94, 95, 96, 97, 98
To address the SIE, we present the generalized r2SCANY@r2SCANX approach, which builds upon r2SCAN. In r2SCANY@r2SCANX, the SIE is addressed by independently tuning the fraction of exact HF exchange directly in the r2SCAN functional definition used to evaluate the total energy of a given set of orbitals and the r2SCAN functional used to compute these orbitals. Specifically, in r2SCANY@r2SCANX, Y is the percent of HF exact exchange appearing in the XC potential of r2SCAN to correct the functional-driven error. In contrast, in r2SCANY@r2SCANX, X is the percent of the exact Hartree-Fock exchange used in the r2SCAN functional to determine the electronic orbitals and potentially correct for the density-driven error. By independently varying the Xand Yparameters in r2SCANY@r2SCANX, the proposed method allows for a more systematic correction of the functional- and density-driven errors inherent to r2SCAN’s formulation, with the aim of improving the accuracy in describing transition metal oxides and potentially other correlated systems.
In r2SCANY@r2SCANX, the XC energy functional ($E_{xc}=E_{xc}^{Y}[n^{X}]$) is defined:
$E_{xc}^{Y}[n^{X}]=\frac{Y}{100}E_{x}^{\mathrm{HF}}[n^{X}]+\left(1-\frac{Y}{100}\right)E_{x}^{\mathrm{r^{2}SCAN}}[n^{X}]+E_{c}^{\mathrm{r^{2}SCAN}}[n^{X}]$ (2)
where, in Eq. 2 $n^{X}$ is the self-consistent electron density obtained from the r2SCANX Hamiltonian, which includes Xexact exchange. Therefore, the r2SCANY@r2SCANX definition incorporates different fractions of exact exchange in both the functional definition and the charge density, ensuring a more systematic correction of SIE. Note, that replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have minimal or no effect on the satisfaction of exact DFT constraints of this meta-GGA.
Using the definition in Eq. 2 and selecting specific values of X and Y, we propose different r2SCANY@r2SCANX XC functionals (Fig. 1). In Fig. 1, the SCF column indicates the r2SCANX part of the functional, whereas the non-self-consistent (non-SCF) is the r2SCANY part. In the SCF part, functionals, such as LDSA-PW92,17 the GGA PBE,18 the meta-GGAs r2SCAN19, 99 (plain or hybrid), or LAK were used.22
Figure 1: Schematically connecting the self-consistent and non-self-consistent approaches required in the r2SCANY@r2SCANX methods. Self-consistent functionals used in this work are LSDA, PBE, r2SCAN, r2SCAN10, and LAK. Non-self-consistent hybrid functionals, including a fraction of exact HF exchange, such as r2SCAN@HF, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, r2SCAN10@r2SCAN50, r2SCAN@LAK, r2SCAN10@LAK, require a self-consistent step to generate orbitals that are evaluated non-self-consistently.
To clarify the exchange and correlation functional nomenclature exposed in Eq. 2 and Fig. 1, we explain, for example, r2SCAN10@r2SCAN50, where the r2SCAN50 functional is used to generate monoelectronic orbitals self-consistently, and these orbitals are then used in combination with a r2SCAN10 functional to compute the total energies non-self-consistently. Similarly, for r2SCAN10@r2SCAN, r2SCAN is used to generate orbitals, and r2SCAN10 is used to calculate the total energy non-self-consistently. Here, specific values of X and Y are selected by error minimization in oxidation energies of MiOjs, which is amply justified. The following sections will return to the choice of X and Y values.
II.2 r2SCAN Predictions of Transition-metal Oxides Properties
[…]
Table 1 summarizes the experimentally predicted lattice parameters, magnetic moments, and band gaps of all materials using r2SCAN and r2SCAN+U (with U parameters calibrated in Ref. 39). All calculations reported in Table 1 used r2SCAN geometry-optimized structures, including the relaxation of atomic positions, volumes, and lattice shapes. Note, for the r2SCAN+U data set, the same functional was used for geometry relaxation (atomic positions, volumes, and shapes).
Table 1: r2SCAN and r2SCAN+U predicted lattice parameters, on-site magnetic moments (in $\mu$B), and band gaps (in eV) of MiOjs. Experimental quantities are marked as Exp. All calculations employ the experimentally-determined magnetic orderings (M.O.) and referenced in the magnetic moment (M.M.) column. Antiferro magnetic (AFM), ferrimagnetic (FEM), ferromagnetic (FM), and non-magnetic orderings (NM) are indicated.
System Method a b c $\pmb{\alpha}$ $\pmb{\beta}$ $\pmb{\gamma}$ M.M. M.O. Band Gap
TiO2 ($P4_{2}/mnm$)100 Exp. 4.59 4.59 2.96 90 90 90 0.00 NM 3.0118
r2SCAN 4.60 4.60 2.96 90 90 90 0.00 NM 2.24
r2SCAN+U 4.62 4.62 2.99 90 90 90 0.00 NM 2.51
Ti2O3 ($R\overline{3}c$)101 Exp. 5.43 5.43 5.43 56.57 56.57 56.57 0.03/0.2101, 119 AFM 0.20120
r2SCAN 5.46 5.46 5.46 57.75 55.75 55.75 0.00 AFM 0.00
r2SCAN+U 5.48 5.48 5.48 57.44 57.44 57.44 0.91 AFM 1.30
VO ($Fm\overline{3}m$)102 Exp. 2.88 2.88 4.99 73.22 90 120 N/A AFM N/A
r2SCAN 3.16 3.16 4.89 71.12 90 120 2.45 AFM 1.66
r2SCAN+U 3.16 3.16 5.00 71.54 90 119.99 2.55 AFM 2.35
V2O3 ($I2/a$)103 Exp. 7.25 5.00 5.55 90 96.75 90 1.2/2.37121, 122 AFM 0.20123
r2SCAN 7.28 5.00 5.51 90 97.50 90 1.70 AFM 0.00
r2SCAN+U 7.28 5.08 5.56 90 96.42 90 1.80 AFM 0.68
V2O5 ($Pmmn$)104 Exp. 11.51 3.56 4.37 90 90 90 0.00 NM 2.5124
r2SCAN 11.59 3.55 4.25 90 90 90 0.00 NM 2.04
r2SCAN+U 11.59 3.56 4.25 90 90 90 0.00 NM 2.14
Cr2O3 ($R\overline{3}c$)105 Exp. 4.95 4.95 13.60 90 90 120 2.76125 AFM 3.2126
r2SCAN 4.94 4.94 13.62 90 90 120 2.58 AFM 2.58
CrO3 ($C2cm$)106 Exp. 4.79 8.56 5.74 90 90 90 0.00 NM 3.8127
r2SCAN 4.86 8.25 5.70 90 90 87.91 0.00 NM 2.30
CrO2 ($P4_{2}/mnm$)107 Exp. 4.42 4.42 2.92 90 90 90 2.00128 FM 0.00128
r2SCAN 4.40 4.40 2.91 90 90 90 2.06 FM 0.00
MnO ($Fm\overline{3}m$)108 Exp. 3.14 3.14 6.29 60 60 60 4.58129 AFM 3.6/3.8130
r2SCAN 3.14 3.14 6.17 59.36 59.36 59.98 4.30 AFM 1.74
r2SCAN+U 3.15 3.15 6.21 59.50 59.50 59.99 4.42 AFM 2.13
MnO2 ($P4_{2}/mnm$)109 Exp. 4.40 4.40 2.87 90 90 90 2.35131 AFM 0.27/0.3132, 133
r2SCAN 4.38 4.38 2.86 90 90 90 2.62 AFM 0.39
r2SCAN+U 4.39 4.39 2.88 90 90 90 2.77 AFM 0.74
Mn3O4 ($I4_{1}/amd$)110 Exp. 5.76 5.76 6.24 117.52 117.52 90 4.34, 3.25/3.6134 FEM 2.3-2.5135
r2SCAN 5.72 5.72 6.22 117.39 117.39 90 4.19, 3.51 FEM 0.96
r2SCAN+U 5.76 5.76 6.24 117.50 117.50 90 4.36, 3.64 FEM 1.39
FeO ($Fm\overline{3}m$)112 Exp. 6.08 6.08 6.08 60 60 60 3.32/4.2136, 137 AFM 2.20138
r2SCAN 5.87 6.13 5.96 62.31 60.48 61.39 3.42 AFM 0.43
r2SCAN+U 6.11 6.10 6.10 61.04 59.91 59.95 3.54 AFM 1.58
Fe2O3 ($R\overline{3}c$)111 Exp. 5.03 5.03 13.76 90 90 120 4.9139 AFM 2.20140
r2SCAN 5.00 5.00 13.74 90 90 120 3.86 AFM 1.52
r2SCAN+U 5.04 5.04 13.75 90 90 120 4.12 AFM 1.50
Fe3O4 ($Fd\overline{3}m$)113 Exp. 8.39 8.39 8.39 90 90 90 4.44, 4.1113 FEM 0.14141
r2SCAN 8.34 8.34 8.34 90 90 90 3.73; 3.70 FEM 0.00
r2SCAN+U 8.44 8.47 8.37 90.01 90.28 90.03 4.12; 3.58 FEM 0.23
CoO ($Fm\overline{3}m$)141 Exp. 3.01 3.01 6.03 60 60 60 3.35/3.8136, 142 AFM 2.40143, 144
r2SCAN 2.99 2.99 5.96 59.92 59.92 60 2.54 AFM 0.85
r2SCAN+U 3.01 3.01 5.97 59.78 59.78 60 2.62 AFM 2.12
Co3O4 ($Fd\overline{3}m$)114 Exp. 8.07 8.07 8.07 90 90 90 3.02145 AFM 1.60144
r2SCAN 8.03 8.03 8.03 90 90 90 2.45 AFM 1.12
r2SCAN+U 8.06 8.06 8.06 90 90 90 2.57 AFM 1.94
CuO ($C2/c$)115 Exp. 9.36 3.42 10.27 90 99.78 90 0.68146 AFM 1.40147
r2SCAN 8.49 3.90 10.27 90 95.25 90 0.56 AFM 0.69
Cu2O ($Pn\overline{3}m$)116 Exp. 4.27 4.27 4.27 90 90 90 0.00 NM 2.17/2.424, 147
r2SCAN 4.24 4.24 4.24 90 90 90 0.00 NM 0.81
Previous studies have shown that r2SCAN and SCAN geometries agree well with experimental data.23, 25, 148, 26, 38, 39 Data in Table 1 confirms the accuracy of r2SCAN in predicting the lattice parameters and magnetic configurations of MiOjs and is consistent with Ref. 39. Exceptions to this trend are VO ($\sim$8.9%) and CuO ($\sim$14%), which show some deviations in their lattice constants from experimental reports.102, 115 Mn3O4 is known to exhibit complex magnetism,134 which was approximated by setting a collinear magnetic configuration that best approximates Mn3O4 experimentally observed magnetic ground state. Experimentally, Ti2O3 ($\sim$0.2 eV), V2O3 ($\sim$0.2 eV), and Fe3O4 ($\sim$0.14 eV) have small experimental band gaps, respectively, but are predicted to be metallic with r2SCAN and consistent with previous results in Ref. 39.
II.3 Benchmarking Oxidation Enthalpies of MiOjs with r2SCANY@r2SCANX Functionals
The XC functionals of Fig. 1 were used to assess the oxidation enthalpies of several binary oxides, MiOjs (M = Ti, V, Cr, Mn, Fe, Co, and Cu) as defined for $z>x$ in Eq. 3.
$\frac{2}{z-x}\mathrm{MO_{x}}+\mathrm{O_{2}}\rightarrow\frac{2}{z-x}\mathrm{MO_{z}}$ (3)
Following Eq. 3, and using various r2SCANY@r2SCANX functional definitions, we computed all possible oxidation enthalpy reactions using the MiOjs in Table 1 and Eq. 4, resulting in 15 distinct reactions. The negative chemical energy change in the reaction is:
$\mathrm{\Delta}H_{0}=\frac{2}{z-x}E(\mathrm{MO_{z}})^{\mathrm{r^{2}SCANY@r^{2}SCANX}}-\frac{2}{z-x}E(\mathrm{MO_{x}})^{\mathrm{r^{2}SCANY@r^{2}SCANX}}-E(\mathrm{O_{2}})^{\mathrm{r^{2}SCANY@r^{2}SCANX}}$ (4)
where, $E\left(\mathrm{MO_{z}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{MO_{x}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, $E\left(\mathrm{O_{2}}\right)^{\mathrm{r^{2}SCANY@r2SCANX}}$, are the r2SCANY@r2SCANX total energies of the oxidized, reduced MiOjs, and $\mathrm{O_{2}}$ gas, respectively. Our DFT predictions are compared to the experimental oxidation enthalpies (at 298 K, and 1 atm) of $\mathrm{MO_{x}}$ and $\mathrm{MO_{z}}$ (Supplementary Table 1) extracted from experimental reports.149, 150, 151 Thermochemical table provide oxidation energies earnest estimates of the corresponding 95% confidence intervals of MiOjs. In addition, here we approximated the oxidation enthalpies by the DFT total energies by imposing $\Delta H_{0}\approx\Delta E$, thus neglecting the pV contributions, which are expected to be minimal.
Fig. 2 displays the deviation of predicted oxidation enthalpies including common XC functionals, such as LSDA,17 PBE,18 and meta-GGA, including r2SCAN99 and LAK,22 as well as r2SCANY@r2SCANX (r2SCANY@LAK) derivatives. From the benchmarking of 15 predicted oxidation energies with experimental values, we derived the error distributions in Fig. 2b, and the mean absolute errors (MAEs) of Fig. 2c. Note, errors in oxidation energies with all methods were evaluated using r2SCAN geometries. For the r2SCAN+U dataset, the MiOjs geometries were optimized using r2SCAN+U.
Figure 2: Prediction error of MiOj oxidation energies with several XC functionals, including r2SCANY@r2SCANX and r2SCANY@LAK as defined in Fig. 1. (a) Error in oxidation energies of all (15) reactions considered. (b) Violin representation of error distributions. (c) The mean absolute errors. Errors in predicting oxidation energies of all DFT functionals except r2SCAN+U are evaluated at r2SCAN geometries. The mean experimental oxidation energy is –3.82 eV/O2.
As shown in Fig. 2b and Fig. 2c, the errors in predicting oxidation energies for MiOjs using LSDA, PBE, and r2SCAN reduce systematically as the quality of the exchange-correlation functional increases from LSDA to PBE, and to meta-GGA. In this sequence, the number of exact constraints on the exchange-correlation energy that can be satisfied increases from 8 to 11 to 17. Predictions with these XC functionals are affected to different extents by the pernicious SIE.15 Therefore, this study will focus on r2SCAN and its hybrids r2SCANY@r2SCANX (Fig. 1).
In Fig. 2c, r2SCAN exhibits a mean absolute error (MAE) of $\sim$1.19 eV/O2 and a root mean squared error (RMSE) of $\sim$1.42 eV/O2 in predicting oxidation enthalpies in our dataset. r2SCAN systematically makes oxidation energies too negative across all tested cases. This is largely a functional-driven error. The r2SCAN+U, which implements an ad hoc U correction.29, 39, 152 The U parameters used here are from Ref. 39. When applied to the 15 oxidation reactions, r2SCAN+U yields significantly milder deviation from experimental data, with an MAE of $\sim$0.57 eV/O2 and an RMSE of $\sim$0.66 eV/O2.
A self-consistent global hybrid, such as r2SCAN10, incorporating 10% of exact HF exchange, should partly mitigate the SIE inaccuracies of the r2SCAN meta-GGA, but at the expense of slightly higher computational costs than for GGA-based global hybrids (and considerably higher than the cost of pure meta-GGAs). As expected, the r2SCAN10 functional significantly reduces the errors, yielding an MAE of $\sim$0.73 eV/O2 and an RMSE of $\sim$0.80 eV/O2 (Fig. 2). This represents approximately a 40% reduction in errors compared to r2SCAN.
When applying r2SCAN@HF and utilizing the HF electronic density for the r2SCAN total energy evaluation, the resultant MAE for our dataset is 1.69 eV/O2 and an RMSE of 2.03 eV/O2. This indicates a decline of approximately 40% in accuracy compared to standard r2SCAN. Similarly to r2SCAN@HF, one can reduce the fraction of HF exact exchange to 50% as in r2SCAN@r2SCAN50 while producing a potentially lower density-driven error than r2SCAN.70, 153 This method gives our dataset an MAE of 0.86 eV/O2 (Fig. 2c) and an RMSE of 0.96 eV/O2.
Building upon r2SCAN@HF, a novel r2SCAN derivative proposed here is the r2SCAN10@r2SCAN, which iterates to self-consistency with the less expensive r2SCAN, requires only a single total-energy evaluation (not requiring a complete self-consistent field electronic relaxation), with the more costly global hybrid r2SCAN10. Supposedly, r2SCAN10@r2SCAN corrects functional-driven errors by introducing 10% exact exchange in the functional while using r2SCAN orbitals.
In Fig. 2b and Fig. 2c, r2SCAN10@r2SCAN yields an MAE of $\sim$0.74 eV/O2 and a RMSE of 0.82 eV/O2, which are comparable in magnitude to the hybrid functional r2SCAN10 but can be obtained at a far lower computational cost. Indeed, r2SCAN10@r2SCAN appears sufficient to correct the functional-driven error of r2SCAN (Fig. 2b). We propose a more general approach in the form of r2SCAN10@r2SCAN50 that corrects both functional-driven errors, including 10% exact Hartree-Fock exchange in total energy estimation, and density-driven errors with 50% exact exchange in the density (orbital) generation. Applying r2SCAN10@r2SCAN50 to our dataset results in an MAE of 0.50 eV/O2 and an RMSE of 0.69 eV/O2. Therefore, r2SCAN10@r2SCAN50 features as the most accurate approach among the r2SCANY@r2SCANX proposed here.
The XC functional LAK is expected to predict more accurate band gaps, which are closely related to charge transfer processes in materials. For this reason, LAK is generally believed to yield improved electron densities compared to r2SCAN,22, 154 and hence, speculatively, r2SCAN@LAK and r2SCAN10@LAK should predict oxidation energies in better agreement with experimental data. However, for our dataset, the performance of r2SCAN@LAK (MAE: $\sim$1.21 eV/O2) appears comparable to that of r2SCAN (MAE: 1.19 eV/O2), and similarly, r2SCAN10@LAK (MAE: 0.81 eV/O2) shows no evident improvements over r2SCAN10@r2SCAN (MAE: 0.74 eV/ O2).
It is essential to analyze the type of distribution of oxidation energies in Fig. 2b. In Fig. 2b, oxidation energies with PBE form a largely unimodal distribution with a long tail towards positive error, comprising reactions such as Fe3O4 $\rightarrow$ Fe2O3, Cu2O $\rightarrow$ CuO, Ti2O3 $\rightarrow$ TiO2, and FeO $\rightarrow$ Fe2O3. The negative tail of this distribution is set by the reaction CoO $\rightarrow$ Co3O4. In contrast, r2SCAN oxidation energies follow a bimodal distribution (Fig. 2b), with low and high error peaks. The low (i.e., close to 0) error peak primarily comprises early transition metal, such as Ti, V, Cr, and Cu reactions. The high error peak includes almost all the Mn, Fe, and Co reactions (Fig. 1a). r2SCAN10@r2SCAN data also follows a bimodal distribution with positive and negative error peaks. Here, the positive error peak is entirely characterized by the Cr reactions, and the rest of the reactions fold into the negative error peak.
The distribution for r2SCAN@r2SCAN50 appears largely unimodal with a broad tail in the positive error comprising Cr2O3 $\rightarrow$ CrO2, and Cu2$\rightarrow$ CuO reactions. r2SCAN10@r2SCAN50 also displays a unimodal distribution with the positive part encompassing the Cr2O3 $\rightarrow$ CrO2 reaction, whereas the negative end of the tail gathers Ti2O3 $\rightarrow$ TiO2, and Fe3O4 $\rightarrow$ Fe2O3.
II.4 Achieving Optimal Fractions of Exact Exchange in the r2SCANY@r2SCANX Formulations
We now focus on potential improvements to the r2SCAN functional, as elucidated in Fig. 1 We will explain how optimal X and Y percentages of exact Hartree-Fock exchange are incorporated in the r2SCAN functionals during the non-self-consistent and self-consistent steps. By utilizing all the oxidation energies predicted (Fig. 2) with various r2SCANY@r2SCANX functionals, we identify the optimal combination of X and Y percentages of HF exchange that minimizes their prediction errors in Fig. 3. In Supplementary Fig. 1, Supplementary Fig. 2, and Supplementary Fig. 3 (which also shows the individual reactions), the behavior of errors for individual reactions appears non-trivial; no correlation could be identified for any of the reactions. Analysis of the MAE and RMSE trends from Fig. 3a and Fig. 3b for the self-consistent r2SCANX and non-self-consistent r2SCANY@r2SCAN methods reveals that the optimal fraction of exact exchange lies in the 10–15% range.
Figure 3: Identification of optimal values of X and Y in r2SCANY@r2SCANX functionals, by locating minima in mean absolute and root mean square errors of oxidation energies of all possible oxidation reactions of MiOjs. Panel a shows that the mean absolute error minimizes at 55.1% exact exchange in orbitals and 10% exact exchange in the functional, forming r2SCAN10@r2SCAN55, which produces an error of $\sim$0.46 eV/O2. Panel b shows that the root mean square error minimizes, with an error value of $\sim$0.66 eV at 44.8% exact exchange in orbital and 8.7% exact exchange in functional, thus r2SCAN9@r2SCAN45. See the method section for details on the interpolation scheme to coarse-grain values of X and Y.
Looking at the x-axis of Fig. 3a and Fig. 3b, for r2SCAN@r2SCANX, the minimum error (both MAE and RMSE) is observed in the range $\sim$55–58exact Hartree-Fock exchange, indicating a significantly higher requirement for exact exchange in the Hamiltonian used for the orbital generation when the underlying functional is r2SCAN. From the global minima in Fig. 3a and Fig. 3b, the r2SCANY@r2SCANX approach achieves its lowest MAE and RMSE when Y –the HF fraction in the functional definition– is in the range of $\sim$8–10%, and X (in the HF fraction in orbitals) lies between $\sim$44–56%.
Although the fraction of exact exchange minimizing the MAE and RMSE differs across methods, a general trend emerges: the error is minimized when approximately 10% exact Hartree-Fock exchange is used in the functional definition, and $\sim$50is included in the orbital generation. These values provide a balanced correction of the SIEs intrinsic to the density functional approximation.
II.5 Effects of Exact-exchange Fractions X and Y on the Binding Energy of the Oxygen Molecule
Figure 4: Error in the negative binding energy of the O2 molecule using various r2SCANY@r2SCANX derivatives. In panel a, the mean absolute error of the same functional definitions as shown in Figure 1. All the calculations are performed using r2SCAN geometries. In panel b, optimal values of X and Y, in various r2SCANY@r2SCANX functionals for O2 binding energy, are demonstrated using the generalized r2SCANY@r2SCANX method. See the method section for details on the interpolation scheme to coarse-grain values of X and Y in panel b.
Fig. 4 displays the error of O2 binding energies introduced by several XC functionals, including the new r2SCANY@r2SCANX proposed here. Fig. 4a shows that all XC functionals considered here tend to overbind the formation of the O2 molecule, with errors diminishing progressively from LSDA-PW92 (–2.2 eV/O2) $\gg$ GGA PBE ($\sim$–1.0 eV/O2) $\gg$ meta-GGA r2SCAN and LAK (–0.3 — –0.2 eV/O2) $\gtrsim$ r2SCANY@r2SCANX.
Note that zero-point energy corrections are not included in this analysis and represent a constant, monotonic energy shift of $\sim$0.1 eV/O2.155, 156 The inadequacy of LSDA XC functionals, which drastically overbinds O2, had already been noted by Perdew and Zunger157 in agreement with LSDA-PW92 in Fig. 4a. Similar inaccuracies were also identified for more accurate GGA functionals by Blöchl158 and later by Wang et al.29
In both LSDA and GGA, the binding energy of O2 is strongly overestimated. This introduces a non-negligible systematic error in the predicted oxidation energies of MiOjs, which could be corrected ad hoc in Eq. 3. In contrast, both SCAN19 and r2SCAN99 meta-GGA significantly alleviate the O2 overbinding of PBE. As shown in Fig. 4a, the error in binding energy drops from $\sim$–1 eV/O2 with PBE to $\sim$–0.3 eV/O2 with r2SCAN. However, due to this non-negligible error in binding energy with r2SCAN, a systematic error persists in the r2SCAN-predicted oxidation energies.
Therefore, we investigate the accuracy of r2SCANY@r2SCANX functionals in O2 binding energies, particularly the optimal X and Y Hartree-Fock exchange values in defining r2SCANY@r2SCANX. Fig. 4 suggests that O2 binding energies can be accurately predicted to eliminate systematic errors and achieve accurate oxidation energies of MiOjs. As indicated by the contour line in Fig. 4b, setting Y at approximately 10% of the exact HF exchange in the non-self-consistent part, along with any value of Xfor exact exchange in the orbital definition, results in a small error in O2 binding energy ($\sim$–0.03 eV/O2).
For r2SCANY@r2SCANX, in Fig. 4a r2SCAN10, r2SCAN10@r2SCAN, and r2SCAN10@r2SCAN50 further reduce the error in O2 binding energy to –0.031, –0.030, and 0.002 eV/O2, respectively, thereby minimizing inaccuracies in predicting oxidation energies of MiOjs. Additionally, r2SCAN@r2SCAN50, which only deals with the density-driven error, does not seem to improve the O2 binding energy error compared to r2SCAN, indicating that the overbinding in O2 binding energy is almost entirely due to the functional-driven error.
II.6 Effects of Exact-Exchange Fractions X and Y on On-site Magnetic Moments
The strong correlation in MiOjs seems to be captured at least in part by density functional approximations for normal correlation through spin symmetry breaking.24, 159, 160, 161, 162 In this work, we have used experimental or nearly experimental magnetic orders (referenced in Table 1).
The artificial delocalization of d electrons due to SIE of XC functionals affects the predicted on-site magnetic moments in transition metals.46, 47, 51, 67, 163, 164, 165 We investigate the variation in error of predicted magnetic moments in our proposed methods r2SCANY@r2SCANX. The magnetic moments are calculated by integrating the net spin density over the projector augmented wave (PAW) potential spheres of the transition metal atoms.
The calculated on-site magnetic moments of transition metals depend only on the accuracy of electronic orbitals, particularly the fraction of exact exchange used in generating those orbitals. The on-site magnetic moments can be considered independent of the functional or the percentage of exact HF exchange employed in the non-self-consistent step (i.e., Y in r2SCANY@r2SCANX), which only affects the energy evaluation and not the electronic charge density. Therefore, only the X percent of exact exchange in defining the r2SCANY@r2SCANX functional is relevant.
Fig. 5a and Fig. 5b show the mean percent and absolute errors in magnetic moments of MiOjs.
Figure 5: Prediction errors of magnetic moments of MiOjs. Panels a and b show percent and non-relative errors in magnetic moments vs. X % exact HF exchange in r2SCANX. Panels c and d show distributions and mean absolute errors of magnetic moments for various XC functionals of interest in this work, such as LSDA-PW91, PBE, r2SCAN, r2SCAN+U, r2SCAN10, r2SCAN50, and HF. The mean experimental magnetic moment is $\sim$2.21 $\mu$B. See Supplementary Table 2 for numerical values of computed magnetic moments.
Increasing the percentage X of Hartree-Fock exchange increases the on-site magnetic moment associated with d electrons. For example, in Fe2O3, the average magnetic moment on Fe centers increases from 3.86 $\mu_{B}$ with r2SCAN to 4.01 $\mu_{B}$ with r2SCAN10, 4.31 $\mu_{B}$ with r2SCAN50, 4.48 $\mu_{B}$ with r2SCAN100, and 4.51 $\mu_{B}$ with Hartree-Fock, i.e., X = 100% and no correlation in the functional definition. This trend aligns with the expected behavior of hybrid functionals, enhancing the localization of magnetic moments. Supplementary Table 1 delineates this trend for all compounds investigated here.
From Fig. 5a, the percent MAE and RMSE of magnetic moments are minimized in an interval of HF exact exchange in the r2SCANX functional definition. There is no significant improvement in percent MAE for HF fractions larger than 45%, whereas percent RMSE increases progressively. Conversely, as shown in Fig. 5b, the absolute magnitude of errors in magnetic moments decreases gradually as the fraction of exact HF exchange increases in defining the orbitals.
Fig. 5c and Fig. 5d demonstrate that, starting from r2SCAN magnetic moments, r2SCAN50 reduces the MAE by 15%, and HF reduces it by 20%. The MAE is especially sensitive to MiOjs with larger magnetic moments (e.g., MnO, FeO, Fe2O3, and Fe3O4 in Table 1, where relatively small fractional moment changes lead to larger absolute differences. However, the fact that the distribution center in Fig. 5c falls near the zero-error line indicates that r2SCAN50 yields a better electron density in this sense.
Finally, we comment on the performance of the meta-GGA LAK in predicting the on-site magnetic moments in MiOjs. Fig. 5c and Fig. 5d show no noticeable improvement of LAK magnetic moments compared to r2SCAN.
II.7 Effects of Exact-exchange Fractions X and Y on the Prediction of Fundamental Band Gaps
Band gaps of materials are typically affected by SIE, with band gaps often underestimated by LSDA, GGA, and meta-GGA functionals.25, 26, 38, 39, 98, 166, 167, 168 We investigate the error in predicted band gaps using our r2SCANY@r2SCANX method. The fundamental band gap is the difference between the lowest unoccupied and the highest occupied orbital energies. Our orbital energies are expectation values of the r2SCANY one-electron generalized Kohn-Sham one-electron Hamiltonian using the r2SCANX orbitals. Y=100% includes the full Fock operator and strongly overestimates the band gaps.
Figure 6: Error in the fundamental band gap of MiOjs. Panels a and b display the mean absolute and root mean square errors in band gaps for the generalized r2SCANY@r2SCANX functionals. Likewise, panels c and d show the distribution and mean absolute errors in band gap for various r2SCANY@r2SCANX. See the method section for details on the interpolation scheme adopted to coarse-grain the values of X and Y of panels a and b. Mean band gap of MiOj is $\sim$1.87 eV. See Supplementary Table 3 and Supplementary Table 4 for numerical values of computed band gaps.
Fig. 6a and Fig. 6b show the MAE and the RMSE in predicted band gaps using our generalized r2SCANY@r2SCANX method. Predicted band gaps with the r2SCANY@r2SCANX method depend on both the Xof exact HF exchange in the functional for the orbitals and the Yof exact HF exchange in the functional used in the non-self-consistent step to evaluate the orbital energies. For MiOjs (Table 1), Fig. 6a and Fig. 6b show that the MAE minimizes at $\sim$5.0exact HF exchange is used in the orbital-generation (i.e., r2SCANX), and around 8% exact HF exchange in the functional used in the non-self-consistent step (r2SCANY). Similarly, the RMSE is minimized at $\sim$12.5exact HF exchange for both components. Indeed, predicted band gaps attain an optimal value for small percentages of X and Y ($\sim$12.5 %) and increase rapidly beyond these percentages.
Fig. 6c shows the distribution of errors, and Fig. 6d shows the mean absolute error in predicted band gaps with XC functionals of Fig. 1. In Fig. 6b the MAE in band gaps systematically decreases from $\sim$1.33 eV for LSDA (PW91) > $\sim$1.24 eV for GGA(PBE) $\gg$ $\sim$0.84 eV in r2SCAN, and > $\sim$0.67 eV for LAK. Moving from LSDA (PW91) to GGA (PBE) to meta-GGA (r2SCAN and LAK), the number of exact constraints satisfied by the DFA increases progressively.
Turning our attention to r2SCANY@r2SCANX-type functionals, r2SCAN10, r2SCAN10@r2SCAN, and r2SCAN10@LAK have similar MAE errors of $\sim$0.50—0.59 eV and show noticeable improvement over r2SCAN. However, r2SCAN@r2SCAN50 and r2SCAN10@r2SCAN50, which performed well for oxidation energies, performed poorly in predicting band gaps, with substantial MAEs of 1.05 eV and 0.76 eV, respectively. In r2SCAN@r2SCAN50 this increase in error primarily originates from 3 closed d-shell oxides TiO2, V2O5, and CrO3 (Supplementary Fig. 4). Notably, r2SCAN+U with a MAE error of 0.71 eV is outperformed by r2SCAN10@r2SCAN in predicting band gaps.
III Discussion
Standard DFAs suffer from significant SIEs, especially when applied to strongly correlated open-shell transition metal oxides (MiOjs). SIE leads to inaccuracies in predicting many essential properties of MiOjs, including their structural parameters, electronic and magnetic structures, and thermochemical data.
Hybrid XC functionals mix a fraction of exact HF exchange with DFA to improve the accuracy of the electronic structure, thereby correcting functional-driven errors, particularly the SIE. The amount of HF exchange remains a tunable parameter. Strategies proposed to tune the amount of HF exchange in the DFA include: fitting molecular heat of formations to experimental data (Pople’s G1 dataset), for example, the B3PW91 and the B3LYP.66, 169 Appropriate amounts of HF can also be derived from first-principles perturbative theory calculations,69 as is the case for 25% HF exchange in PBE0 and HSE06.70, 71 Galli et al.63 demonstrated that the exact exchange amount can be inferred directly considering the material dielectric constant. Scuseria et al.,72, 73, 74, 75 proposed that the amount of HF can be determined as a function of the electron density at each point in space. We have shown that MiOjs’ oxidation energies generally improve climbing the ladder from LSDA to PBE to r2SCAN, but further improvement is needed.
Unsurprisingly, we have shown that admixing exact HF exchange with r2SCAN achieves better electronic, magnetic, and oxidation energies of MiOjs. For example, r2SCAN10 (with 10% HF exchange) reduced the error in oxidation energies by $\sim$40and by $\sim$30for band gaps of MiOjs compared to r2SCAN. r2SCAN10 also reduces overbinding in O2, which is important for the accurate prediction of oxidation energies of MiOjs.
Addressing Density and Functional Driven Errors Altogether: HF also provides self-interaction-free, albeit uncorrelated, electronic charge densities. Electronic charge densities from HF or hybrid functionals can correct density-driven errors. We have demonstrated that HF (or hybrid functional) electronic charge densities can correct r2SCAN (or its hybrids) total energies for MiOjs by proposing a generalized r2SCANY@r2SCANX method. In r2SCAN@r2SCANX, we mix Yexact HF exchange in the DFA and set the fraction of exact HF exchange used to evaluate the charge density to X%. We have proposed optimal combinations of HF exchange in the charge density and functional definitions, such as r2SCAN10, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, and r2SCAN10@r2SCAN50, which reduce the mean absolute (and relative) errors of predicted electronic and magnetic properties, and oxidation energies of MiOjs.
We have demonstrated that r2SCANY@r2SCANX can match or outperform the parametrized r2SCAN+U. For example, in predicting the oxidation energies of MiOjs, the r2SCAN10@r2SCAN50 performed the best with a mean absolute error of $\sim$0.50 eV/O2, which is lower compared to the r2SCAN+U 0.57 eV/O2. This can be rationalized as independently correcting the functional-driven error with 10% exact HF exchange and the density-driven error with 50% exact HF exchange.
Kulik and collaborators demonstrated that the ability of DFT+U to localize electrons on transition metals can vary significantly from that of hybrid functionals; these latter tend to localize the minority spin density (of the transition metal) away from the metal, and towards oxygen atoms in transition metal complexes.170, 171 We observe the same behavior in MiOjs. In agreement with Kulik et al.,170, 171 we also show that a higher fraction of HF exchange in the XC functional increases the on-site magnetic moments on transition metal species in MiOjs, which is rationalized by the rise of majority spin in the d-manifolds. Considering these observations, we can qualitatively state that an increasing amount of HF in the DFA increases the ionic character of MiOjs. We quantify these behaviors for Fe2O3 Supplementary Fig. 5 by change in minority- and majority-spin electron numbers. In addition, Supplementary Fig. 6 shows a progressive increase in the electron number on the oxygen sites with increasing X in r2SCANX. This is preceded by a similar increase from LSDA to PBE to r2SCAN. The electron number on the transition-atom site decreases with increasing X. Exceptions to this trend are MiOjs, where strong charge disproportionation of the transition metal is favored, for example, Fe3O4 (Fe3+ and Fe2+), Mn3O4 (Mn3+ and Mn4+). As for the magnetic moments, the electron number on an atom is estimated by integrating the electron density over the PAW sphere.
In addition, while investigating transition metal complexes, Ref. 171 concluded that DFT+U charge densities might differ substantially from hybrid functionals. To test this observation, we compare r2SCAN@r2SCAN+U MiOjs band gaps and oxidation energies (Supplementary Fig. 7) to r2SCAN and r2SCAN10@r2SCAN50 results. We demonstrate that r2SCAN@r2SCAN+U oxidation energies and band gaps are comparable to r2SCAN.
In contrast, when using HF densities with r2SCAN, as in r2SCAN@HF, we have shown that MiOj oxidation energies worsen compared to r2SCAN. MAEs increased from 1.19 eV/O2 at r2SCAN to 1.69 eV/O2 at r2SCAN@HF. But, we demonstrated that the MAE in MiOj oxidation energies can be halved (w.r.t r2SCAN@HF) to $\sim$0.86 eV/O2 when implementing hybrid electronic charge densities of r2SCAN admixed with 50% HF exchange as in r2SCAN@r2SCAN50.
The widely accepted standard for representing uncertainties in experimentally obtained oxidation enthalpies and other thermochemical data requires estimates of the 95% confidence intervals, which all thermochemical tables universally follow.149, 150, 151 Ruscic has proposed that the MAE of data might underestimate the suggested thermochemical uncertainty by a factor as low as 3.5$\times$, thereby invalidating numerous assertions of achieving ”chemical accuracy.”172 For this reason, while benchmarking the accuracy of r2SCANY@r2SCANX, we have also examined the impact of the root mean square error.
Reproducing errorless O2 binding energies is essential for accurate predictions of oxidation energies of MiOjs. LSDA and PBE drastically overbind The O2 molecule with large negative binding energies,29, 157, 158 $\sim$–2.2 eV/O2 and $\sim$–1 eV/O2, respectively, resulting in a systematic shift in predicted oxidation energies. With r2SCAN this negative error drops in magnitude to a sizeable value of $\sim$–0.3 eV/O2. Applying a constant shift (of $\sim$0.3 eV/O2) to correct the spurious O2 overbinding in r2SCAN would further underestimate MiOj oxidation energies (see Fig. 2a), further exacerbating the error in r2SCAN predictions. Previously, similar corrections have been applied to PBE O2 and MiOj oxidation energies.29, 30 Notably, we have demonstrated that r2SCAN10@r2SCAN, r2SCAN10, and r2SCAN10@r2SCAN50 decrease the error in the O2 binding energy to values $\leq$ 0.03 eV/O2 in magnitude, and without requiring any fitting procedures.29, 30
Since the errors of O2 binding energy are minimal with r2SCANY@r2SCANX methods, errors in the oxidation energies of MiOjs must reside in the prediction of the reduced and oxidized oxides. This is a bimodal distribution of r2SCAN oxidation energies (Fig. 2b), consisting of a low error peak for early transition metals, i.e., Ti, V, Cr, reactions, and higher errors for late transition metals, i.e., Mn, Fe, Co reactions. Notably, early transition metals (Ti, V, and Cr) involve a small number of d-electrons in their electronic configurations, probably reducing correlation effects, which become more prominent for late transition metals (Mn, Fe, and Co) with more electrons in d-shells. In addition, redox reactions involving exchanges between transition metals’ closed d-shell configurations to open d-shell configurations or vice versa, such as Ti4+(d0) $\rightarrow$ Ti3+ (d1), or Cu+ (d10) $\rightarrow$ Cu2+ (d9), appear to be in better agreement with experimental data. However, milder correlation effects in closed-shell transition metals make r2SCAN sufficient to describe the electronic structure of their oxides (for example, TiO2, V2O5, etc.).
From these observations, we expect that the electron density by r2SCAN may be too delocalized to capture localization effects in late transition metals, i.e., Mn, Fe, Co, containing a larger number of d electrons. Better electronic charge densities as provided by r2SCAN@r2SCAN50, correct the electron number in the d manifolds, recovering a unimodal distribution (Fig. 2b), with a substantial decrease of oxidation energies errors for MiOjs with M = Mn, Fe, and Co. In addition, due to the functional-driven error, the average error r2SCAN@r2SCAN50 is non-zero, resulting in an offset distribution around the zero-error line. However, in r2SCAN10@r2SCAN50 the functional-driven error reduces, and the unimodal distribution centers near the zero-error line.
Magnetic moments are also typically impacted by SIE and are underestimated due to the overdelocalization of d(f) electrons on transition metal centers by GGA and meta-GGA functionals. As we climb the Jacob’s ladder173 of the XC functionals from LSDA (PW91) to GGA, and to meta-GGAs, the error in magnetic moments decreases accordingly, as observed from the functional MAE distributions of Fig. 5a and Fig. 5b. We observed that the error in magnetic moments drops slowly as the percentage of exact HF exchange in the orbital increases. Compared to r2SCAN, for r2SCAN50, the error in magnetic moment reduces by $\sim$15%, and for HF, the error reduces by $\sim$20%. In contrast, r2SCAN+U minimizes the error in magnetic moments much more mildly, $\sim$5%. Note that the variance in experimentally reported magnetic moments can be as high as +1.00 $\mu_{B}$. For example, magnetic moments in V3+ in V2O3 are reported to vary between 1.2 to 2.37 $\mu_{B}$ (Table 1),121, 122 which corresponds to variations larger than 100%. Despite this considerable uncertainty in experimentally reported magnetic moments, we observed a noticeable decrease in MAE of MiOjs predicted magnetic moments from r2SCAN to r2SCAN50.
Band gaps of MiOjs are typically underestimated by standard DFT predictions, with errors rapidly decreasing when using LSDA (PW91) $\gg$ GGA (PBE) > meta-GGAs (r2SCAN and LAK) > Hybrid meta-GGA (r2SCANX), see Fig. 6c and Fig. 6d. Using r2SCANY@r2SCANX, we demonstrated that errors in band gaps are minimized with approximately 5-10% exact exchange in the functional and the orbital definitions, which differs from the optimal HF fractions minimizing oxidation energy errors of MiOjs. We have shown that band gaps predicted by r2SCAN10@r2SCAN are more accurate than r2SCAN+U, with MAEs of 0.53 eV/O2 and 0.71 eV/O2, respectively.
Materials possess many degrees of freedom embedded in their crystal structures. Deformation of octahedra (and cation polyhedra), octahedral tilting, and Jahn-Teller distortions all decrease the crystal symmetry in materials.24, 161 Additionally, specific magnetic orderings of transition metals in MiOjs, typically determined by neutron scattering and magnetic spectroscopies, further reduce crystal symmetries. On the one hand, standard structural techniques, such as X-ray and neutron diffraction, generally show less sensitivity to such distortions and defects. As such, reported crystal structures of materials may be inherently more symmetric than they appear. On the other hand, band-gap measurements are sensitive to symmetry-breaking motifs, defects in materials, and charge and magnetic orderings. Zunger and collaborators demonstrated that DFT of overly symmetrized MiOjs and other materials tends to close band gaps, predicting “false metals”.24, 159, 160, 161
Starting from the highly symmetrized experimental structures, further optimized with r2SCAN, our band-gap predictions of Ti2O3,101 V2O3,103 and Fe3O4113 appear metallic (Table 1). Upon removing all possible symmetry layers (perturbing atomic positions from high-symmetry sites of large supercell models, imposing ground state magnetic orderings, removing intrinsic symmetry of wavefunctions, orbitals, and time reversal), Ti2O3, V2O3, and Fe3O4 remain metallic. However, the experimentally reported band gaps for Ti2O3, V2O3, and Fe3O4 are smaller than or equal to $\sim$0.2 eV (Table 1), which makes band-gap opening unlikely, even with substantial symmetry breaking as demonstrated by our calculations. We found no lattice symmetry breaking for Ti2O3 and V2O3, while we observed lattice symmetry breaking for Fe3O4. However, there can be multiple symmetry breakings,24, 159 and we may not have found the one with the lowest energy.
Computational Efficiency Considerations of r2SCANY@r2SCANX approaches: It is relevant to comment on the computational costs faced by the r2SCANY@r2SCANX approach. Excluding the computational time required to obtain the initial set of r2SCAN orbitals self-consistently (which are used directly in the r2SCAN10@r2SCAN calculation and as a starting guess for r2SCANX calculation), the non-self-consistent r2SCAN10@r2SCAN calculations are approximately n times faster than the corresponding SCF global hybrid r2SCAN10 calculations. Here, n represents the number of electronic SCF iterations typically required for a global hybrid r2SCANX convergence, generally ranging from 10 to 300, depending on system complexity and total energy convergence criteria.
IV Conclusions
We have found that prediction accuracy improves for the transition metal oxides as more exact constraints and appropriate non-bonded norms are satisfied by approximating the density functional for the exchange-correlation energy. The mean absolute error (MAE) of oxidation energies is reduced from $\sim$2.6 eV/O2 with LSDA to $\sim$1.7 eV/O2 with PBE GGA to $\sim$1.2 eV/O2 with r2SCAN meta-GGA. r2SCAN+U, with a material-dependent empirical parameter, reaches an MAE of $\sim$0.5 eV/O2 at a modest cost increase. We have found that we can match or slightly exceed that accuracy without a material-dependent parameter, in r2SCAN@r2SCAN50. In a way, this provides insight into the error, although at a greater cost.
In summary, we have developed a novel strategy termed r2SCANY@r2SCANX to mitigate the pernicious self-interaction error of exchange and correlation functionals for the accurate simulations of electronic, magnetic, and thermochemical properties of transition metal oxides. r2SCANY@r2SCANX utilizes different fractions of exact exchange to define both the energy and the density simultaneously, addressing functional-driven and density-driven inaccuracies.
We identified a dependence of the X and Y optimal percentages of exact Hartree–Fock exchange in r2SCANY@r2SCANX justified by their performance on the oxidation energies in the MiOj dataset, magnetic moments, and band gaps. We demonstrated that predicted uncertainties are minimized for: oxidation energies of MiOjs by r2SCAN10@r2SCAN50, and band gaps with r2SCAN10@r2SCAN. In the case of magnetic moments, we showed a noticeable reduction of the MAE with r2SCAN50. We have demonstrated that the r2SCANY@r2SCANX improves upon the r2SCAN predictions and even outperforms the DFT(r2SCAN)+U method, which is commonly employed to predict strongly correlated materials. We observed that r²SCAN10@r²SCAN methods are computationally more affordable than hybrid functionals while maintaining comparable accuracy. Further studies should investigate the transferability of r2SCANY@r2SCANX to other classes of strongly correlated systems, such as transition metal fluorides (which are more ionic than MiOjs) or complex polyanion systems, such as phosphates, thiophosphates, silicates, etc., which have more covalent characteristics than MiOjs.
Self-consistent r2SCAN10, with 10% exact exchange, already reduces the MAE of the oxidation energies to $\sim$0.7 eV eV/O2. r2SCAN10 improves upon r2SCAN, not only for transition-metal oxides but also for sp-bonded molecules and materials. r2SCAN10 even gives oxide band gaps slightly better than r2SCAN+U (Fig. 6). Its small exact-exchange fraction (10%) reflects the reduced self-interaction error in r2SCAN compared to other GGAs and LSDA functionals. To lessen the MAE of oxidation energies in MiOjs even further, to 0.5 eV/O2, seems to require r2SCAN10@r2SCAN50, which diminishes the density-driven error of r2SCAN and r2SCAN10. r2SCAN10@r2SCAN50 also slightly improves the accuracy of magnetic moments (Fig. 5). Notably, the computationally efficient r2SCAN10@r2SCAN is almost as accurate as self-consistent r2SCAN10 for the oxidation energies. Fig. 3 suggests that r2SCANY@r2SCAN cannot do better than this for the oxidation energies. r2SCAN10@r2SCAN appears to be a useful functional for both main-group and transition-element molecules and materials.
Despite its excellent performance in other areas,22, 154 the LAK meta-GGA has been found less accurate than r2SCAN for the oxidation energies (Fig. 2) and only slightly more accurate than r2SCAN for the band gaps (Fig. 5) of these transition-metal oxides.
Note that the first, second, and third ionization energies of the 3d transition metal atoms also improve from LSDA to PBE to r2SCAN.174 That improvement is dramatic for the second and third ionization energies, which seem free from strong correlation effects.
The present investigation and other recent works82, 83, 84, 91, 93, 174 suggest that, in many but not all situations with main-group and transition-metal elements, the functional-driven errors of r2SCAN energy differences can far exceed the relatively small density-driven errors. This has the important consequence that significant improvements in the r2SCAN energy differences can be achieved by doing the expensive self-consistent iteration and geometry optimization with the computationally efficient r2SCAN meta-GGA (possibly including an efficient van der Waals dispersion corrections), and then applying a more expensive nonlocal functional (e.g., a hybrid or self-interaction correction) in a single post-self-consistent calculation, as we have done here in r2SCAN10@r2SCAN. All properties studied here improve on average from LSDA to PBE to r2SCAN to r2SCAN10@r2SCAN. Replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have little or no effect on the satisfaction of exact constraints.
V First-Principles Calculation Details
All calculations presented in this work are performed using the DFT formalism, as implemented in the Vienna Ab initio Simulation Package (VASP).175, 176, 177 The PAW potentials describe the core electrons.178, 179 The electrons from 3s, 3p, 3d, and 4s orbitals are explicitly considered for the transition metal atoms. Using a PAW set that treats fewer valence electrons explicitly (3s, 3p, and 3d) significantly impacts MiOj oxidation energies and band gaps (Supplementary Fig. 8). The kinetic energy cutoff for the plane waves was set to 700 eV, and the total energy was converged to 10-6 eV per cell. Various DFT exchange and correlation functionals were used in their collinear spin-polarized implementation. These are the LSDA-PW92,17 the Perdew, Burke, and Ernzerhof (PBE),18 the meta-GGA functionals r2SCAN,99 and LAK.22 Hartree-Fock and global hybrid r2SCANX77 calculations were also performed, with percentages of exact Hartree-Fock exchange discussed in the text of the paper. Ground-state magnetic orderings reported experimentally (see Table 1), often derived from neutron diffraction experiments, were imposed in all simulations of MiOjs.
Geometries (coordinates, volumes, and cell shapes) converged when the forces on all atoms were lower than 0.01 eV/Å. All properties were calculated with r2SCAN geometries unless explicitly mentioned. For structure relaxation, a $\Gamma$-centered Monkhorst-Pack180 grid with a density of 48 k-points per Å for all systems. All other calculations used a k-grid of approximately 700 k-points per Å3.
The global hybrid r2SCANX orbitals were evaluated at the following percentages of exact Hartree-Fock exchange, with X = 0%, 10%, 25%, 50%, 75%, and 100%, respectively. The non-self-consistent field r2SCANY energies were assessed with a fine grid of Y values, from 0% to 100%, with 5% intervals. In r2SCANY@r2SCANX calculations, we started from self-consistently converged r2SCANX orbitals (WAVECAR in VASP), which are in turn used for a non-self-consistent single electronic step (ALGO = EIGENVAL in VASP) with the r2SCANY functional. A linear interpolation scheme was employed to obtain the energies, band gaps, and magnetic moment results at intermediate fractions of exact HF exchanges not explicitly calculated.
Due to the high computational complexity of global hybrids used in this work, all DFT calculations are performed using unit cells that can accommodate the expected experimental magnetic ordering for each transition metal oxide. On-site magnetic moments of each transition metal atom were obtained by integrating the spin density within the projection radius of the PAW potentials. Changes of such radii have a negligible influence on the numerical values of the magnetic moment.
For estimating band gaps with r2SCANY@r2SCANX, we utilize the orbitals obtained self-consistently with r2SCANX, subsequently, and apply the r2SCANY functional non-self-consistently on the same orbitals to obtain the one-electron eigenenergies. These one-electron eigenenergies are then used to calculate the band gap.
Author contributions
P.C. and J.P.P. designed and supervised the project. H.R.G. prepared and performed the simulations, data collection, and data analysis. P.C., J.P.P., J.S., R.Z., and Y.W. contributed to the data analysis. H.R.G. and P.C. wrote the first draft of the manuscript. All authors contributed to the discussion and the final version of this manuscript.
Competing Interests
The authors declare no competing interests.
Additional Information
Supplementary Information. The online version contains supplementary material available at
Data Availability
All the computational data associated with this study, including the input and output files of the simulations, are available on Zenodo at https://doi.org/10.5281/zenodo.15741824.
Acknowledgements.
The Welch Foundation is acknowledged for providing P.C. a Robert A. Welch Professorship at the Texas Center for Superconductivity and the Welch Foundation grant No. E-2227- 20250403. We are grateful for the support of the Research Computing Data Core at the University of Houston for assistance with the calculations carried out in this work. J.P.P. acknowledges support from the National Science Foundation under grant DMR-2426775 and the Department of Energy Office of Science under grant DE-SC001833. J.S. acknowledges support from the National Science Foundation under grant DMR-2042618.
Supplementary Figure 4: Absolute error in band gaps for MiOjs.
Supplementary Figure 5: Changes in electron numbers for minority spin and majority spin on the Fe atom in Fe2O3 ($R\overline{3}c$), as functions of the percentage X of exact exchange in r2SCANX.
Supplementary Figure 6: Changes in electron numbers as functions of the percentage X of exact exchange in r2SCANX.
Supplementary Figure 7: Error in oxidation energy and band gap of MiOjs with r2SCAN, r2SCAN@r2SCANX+U, and r2SCAN@r2SCAN50.
Supplementary Figure 8: Error in oxidation energy and band gap of MiOj with r2SCAN using different PAW potentials, 3s 3p 3d 4s and 3p 3d 4s.
Supplementary Table 1: Experimental oxidation enthalpies (${\Delta\mathrm{H}_{0}}$ in eV/f.u.) of MiOjs.
System Expt. $\pmb{\Delta\mathrm{H}_{0}}$ Ref.
TiO2 ($P4_{2}/mnm$)100 –9.74 151
Ti2O3 ($R\overline{3}c$)101 –15.67 151
VO ($Fm\overline{3}m$)102 –4.45 151
V2O3 ($I2/a$)103 –12.58 151
V2O5 ($Pmmn$)104 –15.97 151
Cr2O3 ($R\overline{3}c$)105 -11.81 149
CrO3 ($C2cm$)106 –6.11 149
CrO2 ($P4_{2}/mnm$)107 –6.20 149
MnO ($Fm\overline{3}m$)108 –3.99 151
MnO2 ($P4_{2}/mnm$)109 –5.39 151
Mn3O4 ($I4_{1}/amd$)110 –14.37 151
FeO ($Fm\overline{3}m$)112 –2.74 150
Fe2O3 ($R\overline{3}c$)111 –8.51 150
Fe3O4 ($Fd\overline{3}m$)113 –11.57 150
CoO ($Fm\overline{3}m$)141 –2.47 149
Co3O4 ($Fd\overline{3}m$)114 –9.23 149
CuO ($C2/c$)115 –1.63 149
Cu2O ($Pn\overline{3}m$)116 –1.75 149
Supplementary Table 2: On-site magnetic moments (in $\mu_{\mathrm{B}}$) of MiOjs calculated with LSDA (PW92), PBE, r2SCAN, and r2SCANX hybrid functionals and compared to experimentally reported values.
System Expt. PW92 PBE LAK r2SCAN r2SCAN+U r2SCAN10 r2SCAN25 r2SCAN50 r2SCAN75 r2SCAN100 HF
TiO2 ($P4_{2}/mnm$)100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Ti2O3 ($R\overline{3}c$)101 0.3-0.2101 0.00 0.00 0.00 0.00 0.81 0.57 0.77 0.85 0.88 0.89 0.91
VO ($Fm\overline{3}m$)102 N/A 2.02 2.22 2.47 2.45 2.55 2.52 2.59 2.66 2.69 2.72 2.76
V2O3 ($I2/a$)103 1.2-2.37121, 122 1.25 1.47 1.74 1.70 1.80 1.78 1.83 1.86 1.87 1.87 1.90
V2O5 ($Pmmn$)104 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Cr2O3 ($R\overline{3}c$)105 2.76125 2.37 2.45 2.61 2.58 2.63 2.67 2.71 2.73 2.74 2.77
CrO3 ($C2cm$)106 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
CrO2 ($P4_{2}/mnm$)107 2.00128 1.89 1.94 2.10 2.06 2.13 2.21 1.99 1.95 1.94 1.98
MnO ($Fm\overline{3}m$)108 4.58129 4.08 4.15 4.35 4.3 4.42 4.36 4.42 4.48 4.52 4.54 4.58
MnO2 ($P4_{2}/mnm$)109 2.35131 2.27 2.42 2.67 2.62 2.77 2.71 2.82 2.93 2.99 3.00 3.06
Mn3O4 ($I4_{1}/amd$)110 4.34, 3.86; 3.98; 4.24; 4.19; 4.35; 4.28; 4.36; 4.43; 4.48; 4.51; 4.54;
3.25–3.64134 3.29 3.35 3.55 3.51 3.65 3.57 3.63 3.69 3.70 3.71 3.75
FeO ($Fm\overline{3}m$)112 3.32–4.2136, 137 3.23 3.30 3.46 3.42 3.54 3.47 3.53 3.59 3.63 3.65 3.67
Fe2O3 ($R\overline{3}c$)111 4.9139 3.24 3.45 3.95 3.86 4.12 4.01 4.16 4.31 4.41 4.47 4.51
Fe3O4 ($Fd\overline{3}m$)113 4.44,4.1113 3.19; 3.36; 3.83; 3.73; 4.02; 3.89; 4.04; 4.22; 4.35; 4.43; 4.46;
3.34; 3.34 3.46; 3.46 3.76; 3.76 3.69;3.69 3.58; 4.12 3.74; 3.77 3.62; 4.13 3.58; 4.32 3.62; 4.42 3.64; 4.49 3.66;4.52
CoO ($Fm\overline{3}m$)141 3.35–3.8136, 142 2.31 2.39 2.57 2.54 2.62 2.61 2.67 2.73 2.76 2.78 2.79
Co3O4 ($Fd\overline{3}m$)114 3.02145 2.07 2.21 2.51 2.45 2.57 2.55 2.62 2.69 2.73 2.75 2.76
CuO ($C2/c$)115 0.68146 0.00 0.33 0.61 0.56 0.65 0.74 0.82 0.87 0.90 0.91
Cu2O ($Pn\overline{3}m$)116 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Supplementary Table 3: Band gaps (in eV) of MiOjs calculated with LSDA (PW92), PBE, r2SCAN, and r2SCANY@r2SCANX and compared to experimentally (Exp.) reported values.
System Exp. PW92 PBE r2SCAN r2SCAN+U r2SCAN10 r2SCAN10@r2SCAN r2SCAN@r2SCAN50 r2SCAN10@r2SCAN50 r2SCAN@HF
TiO2 ($P4_{2}/mnm$)100 3.0118 1.76 1.85 2.24 2.51 3.07 3.41 0.61 1.85 0.00
Ti2O3 ($R\overline{3}c$)101 0.2120 0.00 0.00 0.00 1.52 0.82 0.00 0.21 1.52 0.47
VO ($Fm\overline{3}m$)102 N/A 0.00 0.57 1.66 2.35 3.10 2.99 1.98 3.37 2.41
V2O3 ($I2/a$)103 0.2122 0.00 0.00 0.00 0.68 1.01 0.00 0.00 1.19 0.36
V2O5 ($Pmmn$)104 2.5124 1.60 1.70 2.04 2.14 2.90 3.22 0.42 1.67 0.00
Cr2O3 ($R\overline{3}c$)105 3.2126 1.20 1.64 2.58 4.00 3.93 2.65 4.03 2.42
CrO3 ($C2cm$)106 3.8127 1.97 2.05 2.29 3.12 3.31 1.14 2.26 0.00
CrO2 ($P4_{2}/mnm$)107 0.00128 0.00 0.00 0.00 0.23 0.00 0.00 0.00 0.00
MnO ($Fm\overline{3}m$)108 3.6–3.8130 0.89 1.00 1.74 2.13 2.90 2.87 2.05 3.03 2.21
MnO2 ($P4_{2}/mnm$)109 0.27–0.3132, 133 0.00 0.00 0.39 0.74 1.48 1.68 0.0 0 0.33 0.00
Mn3O4 ($I4_{1}/amd$)110 2.3–2.5135 0.15 0.45 0.96 1.39 2.21 2.38 0.41 1.70 0.00
FeO ($Fm\overline{3}m$)112 2.20138 0.00 0.00 0.43 1.58 1.90 1.72 0.77 2.04 0.99
Fe2O3 ($R\overline{3}c$)111 2.20140 0.41 0 .64 1.52 2.71 2.96 2.95 1.19 2.64 1.23
Fe3O4 ($Fd\overline{3}m$)113 0.14141 0.00 0.00 0.00 0.23 0.13 0.00 0.00 0.00 0.00
CoO ($Fm\overline{3}m$)141 2.40143, 144 0.00 0.00 0.85 2.13 2.67 2.57 1.61 3.26 2.37
Co3O4 ($Fd\overline{3}m$)114 1.60144 0.38 0.68 1.12 1.94 2.88 2.86 0.82 2.37 0.29
CuO ($C2/c$)115 1.40147 0.00 0.00 0.69 2.16 1.80 1.45 2.91 1.86
Cu2O ($Pn\overline{3}m$)116 2.17–2.4147, 24 0.58 0.58 0.81 1.68 1.67 0.59 1.58 0.00
Supplementary Table 4: Band gaps (in eV) of MiOjs calculated, with LAK and r2SCANX@LAK and compared with experimentally (Exp.) reported data.
System Exp. LAK r2SCAN@LAK r2SCAN10@LAK
TiO2 ($P4_{2}/mnm$)100 3.0118 2.38 1.91 3.12
Ti2O3 ($R\overline{3}c$)101 0.2120 0.00 0.00 0.00
VO ($Fm\overline{3}m$)102 N/A 2.06 1.72 3.06
V2O3 ($I2/a$)103 0.2122 0.2 0.0 0.63
V2O5 ($Pmmn$)104 2.5124 2.19 1.69 2.9
Cr2O3 ($R\overline{3}c$)105 3.2126 2.96 2.6 3.97
CrO3 ($C2cm$)106 3.8127 2.35 2.01 3.05
CrO2 ($P4_{2}/mnm$)107 0.00128 0.00 0.00 0.00
MnO ($Fm\overline{3}m$)108 3.6–3.8130 2.20 1.91 2.94
MnO2 ($P4_{2}/mnm$)109 0.27-0.3132, 133 0.61 0.14 1.45
Mn3O4 ($I4_{1}/amd$)110 2.3–2.5135 1.26 1.00 2.25
FeO ($Fm\overline{3}m$)112 2.20138 0.50 0.44 1.71
Fe2O3 ($R\overline{3}c$)111 2.20140 1.99 1.51 2.90
Fe3O4 ($Fd\overline{3}m$)113 0.14141 0.00 0.00 0.00
CoO ($Fm\overline{3}m$)141 2.40143, 144 1.06 0.91 2.58
Co3O4 ($Fd\overline{3}m$)114 1.60144 1.69 1.46 2.86
CuO ($C2/c$)115 1.40147 1.04 0.71 1.93
Cu2O ($Pn\overline{3}m$)116 2.17–2.4147, 24 0.99 0.82 1.70